what is closer to 2/5, 6/21 or 13/29?
1
Expert's answer
2020-10-21T17:26:34-0400

Distance between "\frac{2}{5}" and "\frac{6}{21}" is "d\left(\frac{2}{5},\frac{6}{21}\right)=\bigl|\frac{2}{5}-\frac{6}{21}\bigr|=\bigl|\frac{2}{5}-\frac{2}{7}\bigr|=\bigl|\frac{2\cdot 7-2\cdot 5}{5\cdot 7}\bigr|=\bigl|\frac{4}{5\cdot 7}\bigr|=\frac{4}{5\cdot 7}"

Distance between "\frac{2}{5}" and "\frac{13}{29}" is "d\left(\frac{2}{5},\frac{13}{29}\right)=\bigl|\frac{2}{5}-\frac{13}{29}\bigr|=\bigl|\frac{2\cdot 29-13\cdot 5}{5\cdot 29}\bigr|=\bigl|\frac{-7}{5\cdot 29}\bigr|=\frac{7}{5\cdot 29}"

We have "d\left(\frac{2}{5},\frac{6}{21}\right)=\frac{4}{5\cdot 7}=\frac{4\cdot 29}{5\cdot 7\cdot 29}>\frac{7\cdot 7}{5\cdot 7\cdot 29}=\frac{7}{5\cdot 29}=d\left(\frac{2}{5},\frac{13}{29}\right)" .

So "\frac{13}{29}" is closer to "\frac{2}{5}" than "\frac{6}{21}".

Answer: "\frac{13}{29}"